Simplify. X^p+(1/2)q ×Y^2p-q ÷ (XY^2)^P × √X^q
1
Expert's answer
2020-07-08T17:21:52-0400

I assume your question is: "x^{p+\frac q 2}\times y^{2p-q} \div (xy^2)^p \times (\sqrt {x})^q"

Then the solution is:

"x^{p+\frac q 2}\times y^{2p-q} \div (xy^2)^p \times (\sqrt {x})^q =" "x^{p+\frac q 2}\times y^{2p-q} \times \frac {x^{q/2}}{(xy^2)^p} ="

"=x^{p+\frac q 2+\frac q 2 -p}\times y^{2p-q-2p} =x^q \times y^{-q} = (\frac x y)^q"

Final answer: "(\frac x y)^q"